OpenStudy (anonymous):

Please help will give medal and fan !! Find the counterexample to this statement: If xy = 0, then y must be equal to 0. Select one: a. x(0) = 0 b. (0)(0) = 0 c. (0)y = 0 d. (1)(1) = 1

OpenStudy (whpalmer4):

Hint: 2*0 = 0, and 0*2 = 0

OpenStudy (anonymous):

im still confused. :/

OpenStudy (whpalmer4):

Well, okay. Is \(y=0\) the only way that you can have \(x*y = 0\)?

OpenStudy (whpalmer4):

It seems to me that you can have \(y = 2\) and still have \(x*y = 0\) if \(x = 0\), no?

OpenStudy (anonymous):

umm. yea

OpenStudy (whpalmer4):

Umm, yeah, indeed. So do any of the answer choices suggest a way where \(x*y = 0\) even though \(y \ne 0\)?

OpenStudy (anonymous):

C ?

OpenStudy (whpalmer4):

That would be my choice. It shows a way that \(x*y = 0\) could be true no matter what value \(y\) has. We're looking for a counterexample to the statement that \(x*y = 0\) implies that \(y=0\), and that is one.

OpenStudy (whpalmer4):

It's like saying "all horses have 4 legs. therefore, all animals with 4 legs are horses." to which you respond "a dog has 4 legs, and is not a horse." That's your counterexample.

OpenStudy (anonymous):

lol (: thanks man

OpenStudy (whpalmer4):

You bet! Got any more? :-)

OpenStudy (anonymous):

Sadly, yes. i hate math. ill open a new question

OpenStudy (whpalmer4):

Okay, operators are standing by to field your question :-)